you just received a shipment of six televisions. Two of the televisions are defective If two televisions are randomly selected, compute the probability that both televisions work. What is the probablity at least one of the two televisions does not work? The probability that both televisions work is (Round to three decimal places as needed) The probability that at least one of the two televisions does not work is (Round to throe decimal places as needed Guide uccess Library Options Time Spent h 37m Gm 51m
Answer #1

a)

if both television work then both the television selected from 4 non defective TVs.

Number of ways to select both televisions from 4 non-defective television = 4C2

Total ways to select 2 televisions from 6 available televisions = 6C2

P(Both television work) = 4C2 / 6C2

= 6 / 15

= 0.4

b)

P(At least one television does not work) = 1 - P(Both television work)

= 1 - 0.4

= 0.6